OpenStudy (anonymous):

An appliance store offers a TV set marked at $6,400 less discounts of 10% and 15%. How much does the buyer save?

OpenStudy (anonymous):

A discount of 10% means the new price is 90% of the previous one. A discount of 15% means the new price is 85% of the previous one. So if the original price is P, then the price after the 10% discount is 0.9P. If you apply a 15% discount to that price, the new price is 0.85(0.9P) = 0.765P P is given as 6400 Price is 0.765 * 6400 = $ 4896 Amount saved = 6400 - 4896 = 1504 He saves $ 1504